OpenStudy (bananas1234):

If f(x) varies directly with x2, and f(x) = 75 when x = 5, find the value of f(4). 100 50 48 25

OpenStudy (anonymous):

so... what would your equation be?

OpenStudy (bananas1234):

f(x) = kx

OpenStudy (bananas1234):

@Nnesha Could you help?

jimthompson5910 (jim_thompson5910):

f(x) varies directly with x^2 not just x

jimthompson5910 (jim_thompson5910):

try again @Bananas1234

OpenStudy (bananas1234):

try again for what? the formula?

jimthompson5910 (jim_thompson5910):

yes, it's not f(x) = kx

OpenStudy (bananas1234):

its the direct variation formula right?

jimthompson5910 (jim_thompson5910):

yeah but you are missing something

OpenStudy (bananas1234):

then it is k = f(x)/x to find the constant.

jimthompson5910 (jim_thompson5910):

it's not that either

jimthompson5910 (jim_thompson5910):

f(x) varies directly with x^2 f(x) varies directly with x SQUARED

OpenStudy (bananas1234):

oh sorry i didn't know what you meant

jimthompson5910 (jim_thompson5910):

so what is the true formula here?

OpenStudy (bananas1234):

f(x) = kx^2

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

we're given that x = 5 and f(x) = 75. Use them both to find k

OpenStudy (bananas1234):

k = 75/5? that would = 15. so does k = 15?

jimthompson5910 (jim_thompson5910):

remember, we have x^2 and not just x

OpenStudy (bananas1234):

/5